For a quantitative trait explained by three loci in a additive fashion, if the three locus genotype aabbcc corresponds to a phenotypic value of 10, and the three-locus genotype AABBCC corresponds to a phenotypic value of 40, then what is the phenotypic value corresponding to the three-locus genotype AABbcc?

A. 20

B. 35

C. 30

D. 25

E. 15
Answer #1

Here, we can see that the recessive genotype has phenotypic value 10, this means that this is the minimum value which will be there even when there is no dominant allele.
From the phenotypic value of the dominant genotype AABBCC, we can find the additive value of each dominant allele.

Contribution of one dominant allele = 40-10/ Number of alleles = 30 / 6 = 5

Now in AABbcc, there are three dominant alleles.
So the phenotypic value = 10 + 5 + 5 + 5 = 25.

Correct answer: Option D.